1 Least Elements of Polyhedral Sets and Nonnegative Generalized Inverses

By Debashisha Mishra and Sivakumar K.C..

A classical result due to Cottle and Veinott gives a characterization of the existence of the least element of a specific polyhedral set defined by a matrix, in terms of nonnegativity of a left-inverse of the matrix. In this talk we present extensions of this result to semi-infinite matrices and characterize nonnegativity of certain classes of generalized inverses.